OpenStudy (anonymous):

If the roots of the equation (a²+b²)x²-2(ac+bd)x+(c²+d²)=0 are equal, prove that a/b=c/d.

OpenStudy (anonymous):

this is a math problem but let \[p=a^2+b^2\] \[w=c^2+d^2\] G=ac+bd so you are left with; \[Px^2-2Gx+W=0\] using viete's formula \[x1+x2 = \frac{b}{a}, x1*x2 =\frac{a}{c}\] because of the -b then use the fact that a/b=c/d, and a=(cb)d substitute this in both solutions and you will be done.

OpenStudy (anonymous):

I assumed that P,G and W are real numbers

OpenStudy (anonymous):

In this case; \[x1+x2 = \frac{2G}{P}, x1*x2 =\frac{P}{W}\] because of the -2G